# Improved treatment of relativistic effects in linear augmented plane wave

### v1

@@ Abstract
We examine the influence of the relativistic effects within the linear augmented plane wave method (LAPW) with the potential of general shape for solids and suggest a few ways to account them more accurately: (1) we introduce new radial wave functions based on two actual radial solutions of the Dirac equation for j=l1/2j=l-1/2 and $j=l+1/2$ one-electron states; (2) the canonical LAPW matrix elements for the spherically symmetric component of the potential, assuming non-relativistic radial wave functions, should be corrected; (3) we argue that for a realistic spin-orbit energy splitting of the semicore $6p$-states the spin-orbit interaction constant $\zeta(p)$ should be calculated with the $6p_{3/2}$ radial component; (4) in cases when two $j=l\pm 1/2$ components are occupied (for example for the $6p$ states of actinides) the electron density, associated with the small components of valence electrons, can be taken into the calculation scheme. We demonstrate that the new treatment for the relativistic effects is capable to change the equilibrium lattice constant up to 0.15 Å and the bulk modulus up to 26 GPa. We find that the electron density of valence electrons at the nucleus increases by 2.34.32.3-4.3 times due to the inclusion of small components, which can be essential for precise description of the potential and density close to the nuclear region, important for nuclear spectroscopies. In contrast to the common believe that in plain band structure treatment UO2 is a metal, we show that in the presence of the spin-orbit coupling UO2 has a small gap of forbidden states (0.20.40.2-0.4 eV) at the Fermi level, where the highest occupied and the lowest unoccupied $5f$ bands slightly overlap, as in calculations of the conduction and valence band in solid Ge.
@@ I Introduction
The aim of the present study is to increase the accuracy of the relativistic effects within the full electron full potential LAPW method (FLAPW) method for heavy elements as much as possible, while keeping the general scheme of the method unchanged. For that purpose (1) we introduce new basis functions, obtained from two independent solutions of the Dirac equation; (2) we reconsider the matrix elements of the method, explicitly avoiding the use of hidden non-relativistic relations; (3) we correct the calculation of the spin-orbit (SO) coupling constant for $6p$ semicore states, based on the comparison the energy splittings between $6p_{1/2}$ and $6p_{3/2}$ components. In the present study we do not apply the second variation procedure for the SO-coupling. We use the direct treatment of the SO coupling in the full LAPW basis set thereby avoiding the approximations associated with the second variation step [10]. In this respect it is worth noting that in some relativistic calculations the authors tend to avoid the inclusion of the SO interaction [19, 20, 5], which, as discussed in Sec. II.3 below, can be connected with the overestimated values of SO coupling [6]. Finally, (4) we include the small components $Q$ in the calculations, which play an important role in the neighborhood of the nuclear region. These effects are considered and discussed in Sec. II.1, Sec. II.2, Sec. II.3, and Sec. II.4, correspondingly. In Sec. III we briefly review the results of our calculations for Ac, Th, ThO2 and UO2, and finally in Sec. IVwe discuss main conclusions and findings of our work. In our study we use various variants of the DFT functionals (see Sec. III below), which allows us to test the accuracy of the calculations.
@@ II.3 Calculated spin-orbit coupling constants, special treatment for $6p$-states
From the table 5 we can conclude that the actual energy splittings $\triangle E(p)$ between the $p_{1/2}$ and $p_{3/2}$ states are better approximated by the SO coupling constant $\zeta(p_{3/2})$ calculated using a single radial component $p_{3/2}$, for which $\triangle_{SO}^{*}(p)=3\zeta(p_{3/2})/2$. For example, for Ac we obtain $\triangle_{SO}^{*}(p)=6.09$ eV, for Th $\triangle_{SO}^{*}(p)=7.12$ eV etc. Although the values $\triangle_{SO}^{*}(p)$ are slightly smaller than the real energy differences $\triangle E(p)$, they approximate $\triangle E(p)$ better (i.e. 9119-11% vs 172717-27%) than the KH radial functions $P_{l=1}^{KH}$. Although in this Section we have considered the results with the PBE variant of DFT, the same conclusions can be drawn for other DFT functionals. Therefore, in the following for the LAPW calculations of SOC constants $\zeta(p)$ for the $6p$ semicore band states we will use only $p_{3/2}$ radial component. For the LAPW calculations of the SO couplings $\zeta(d)$, $\zeta(f)$ for the $d$ and $f$ valence states we use the standard averaged radial components, Eq. (3), because, as discussed above, they give good approximations of the energy splittings. Our results with such SO coupling constants are reproduced in Tables 69 below.
@@ III Application to actinides
The technical parameters of numerical calculations were as follows. For fcc Ac and fcc Th, in most cases the number of augmented plane waves was 137 and 274 (with SO), with $R_{MT}K_{j}\sim 10$, where $\vec{K}_{j}$ is the maximal wave vector, and $R_{MT}$ is the radius of the MT-sphere. For CaF2 cubic structures, the basis sets were 307 and 614 (with SO) for ThO2 ($R_{MT}K_{j}\sim 8.7$), 387 and 774 (with SO) for UO2 ($R_{MT}K_{j}\sim 9.5$). For the KH (canonical) basis two sets of $6p$ local functions ($u$ and $\dot{u}$) as described in Ref. [11]. Therefore, for the KH-calculations the total number of basis functions is 143 and 286 (with SO) for Ac and Th, 313 and 626 (with SO) for ThO2, and 393 and 786 for UO2. The following radii of MT-spheres have been used: for Ac 3.46 a.u. for all DFT functionals (i.e. LDA, PBEsol, PBE); for Th 3.05 a.u. (LDA), 3.12 a.u. (PBEsol and PBE); for ThO2 2.25 a.u. for both Th and O for all DFT functionals; for UO2 2.138 a.u. (LDA), 2.168 a.u. (PBEsol, PBE) for both U and O. It is worth noticing that within the chosen DFT functional we used the same $R_{MT}$ for all variants of the different radial functions (i.e. avD, avD(Q), KH with and without SOC). The dependence of the results of the new treatment on the chosen $R_{MT}$ is briefly discussed in Appendix C. The maximal number of $k$-points in the irreducible part (IP) of the Brillouin zone (BZ) for elemental actinides was 1505 ($\sim 70000$ for the whole BZ). For ThO2 and UO2 we have used a set of 240 $k$-points in IP of BZ ($\sim 11500$ for the whole BZ). The maximal value of the LAPW plane-wave expansion and the non-spherical density decomposition was $L_{max}=8$. We have taken into account the finite size of nuclei and used the tetrahedron method for the linear interpolation of energy between $k$-points [29]. The number of radial points inside the MT sphere region was increased to 4000-4200 for actinides and 900-1000 for oxygen. The enlarged number of radial points is analogous to the increase of the quality of the basis set, and this is certainly required for the actinides with 80 core electrons, which is a very large quantity. The calculations with the KH basis sets have been performed with additional localized atomic $p$-basis functions as described in Ref. [11]. New radial basis functions $P_{\ell}^{av}$ (avD) have turned out to work well even without them, and all calculations marked in the Tables as avD or avD(Q) (with and without SO) have been carried out without localized functions.
@@
The results of our calculations are listed in Table 6 for fcc Ac, Table 7 for fcc Th, Table 8 for cubic ThO2, and Table 9 for cubic UO2. ThO2 and UO2 are crystallized in the CaF2 structure. As discussed in the Introduction, most calculations of actinides [2, 4] and their oxides [5, 3, 33, 34] have been performed with the full potential LMTO method. Although nowadays the full potential LAPW (FLAPW) study of actinides [6, 7] and dioxides [35, 36] are also available, the new feature of the present FLAPW calculations is the complete treatment of the SO couplings. It is also worth mentioning that there are several theoretical studies of UO2 with correlation effects (Hubbard repulsion) [37, 38, 39, 40, 41, 42, 43]. Such an approach however lies beyond the scope of the present work, which focuses on the peculiarities of the inclusion of relativistic effects.
@@
It is also worth mentioning that in contrast to Th, our DFT calculations of the fcc structure of Ac appreciably overestimate its lattice constant even for LDA. This however was also noticed e.g. in Ref. [48] ($a=5.503$ Å in LDA) and therefore, the effect should be attributed to the peculiarity of the band structure of this element. Our calculations indicate that this feature becomes more pronounced with increasing quality of the basis set. For example, decreasing the non-spherical components of electron density to $L_{max}=6$ in the avD(Q) basis leads to a smaller lattice constants: 5.536 Å ($\delta a=-0.037$ Å) in LDA, 5.736 Å (0.018-0.018 Å) in PBE, 5.597 Å (0.032-0.032 Å) in PBEsol. Further, decrease of the basis set to only 65 functions results in 5.500 Å (total $\delta a=-0.073$ Å in LDA) 5.608 Å (0.146-0.146 Å) in PBE, 5.539 Å (0.09-0.09 Å) in PBEsol, compare with Table 6. Possibly, some properties of the phonon spectrum and mean square displacements of atoms in solid Ac make the description using poor basis sets more adequate to the experimental data. Owing to its scarcity and radioactivity, the experimental bulk modulus of Ac is unknown. To the best of our knowledge in the literature there is only an estimated (not directly measured) value of 24.5 GPa, listed in Ref. [46]. There is also a theoretical estimation, $B=25.9$ GPa, performed on the basis of a tight-binding analysis [48]. Both values are in good correspondence with our PBE and PBEsol data, Table 6, which represent ab initio theoretical data for bulk moduli of Ac.
@@
Figure 5: The upper panel of the calculated band structure of UO2 with the spin-orbit coupling, PBE calculation. The highest occupied $5f$ electron band is shown by solid line, lowest unoccupied $5f$ bands by dashed lines. The vertical gap with $\triangle E$ from 0.20.40.2-0.4 eV is visible.
@@ IV Conclusions
Within the LAPW method we have presented a few ways to include the relativistic effects more completely and consistently: (1) we have used new radial basis functions $P_{\ell}^{av}$ and $\dot{P}_{\ell}^{av}$, which are obtained by finding the large components $P_{\kappa=\ell}$, $P_{\kappa=-\ell-1}$ of the Dirac solutions independently for the $j=\ell-1/2$ and $j=\ell+1/2$ states and then averaging them explicitly by means of Eq. (3). We have found that new radial basis functions $P_{\ell}^{av}$ bring more stability to the self-consistent-field procedure, performing quite well even without additional $6p$ semicore basis functions used e.g. in Ref. [11], but also without the $6p_{1/2}$ function used in the second variation SO treatment of Refs. [24, 25]; (2) we have corrected the LAPW expressions for $a_{l}^{n}$, $b_{l}^{n}$, Eq. (11a), Eq. (11b), and for the matrix elements in the spherical ($L=0$) component of the total potential, using Eq. (13) for $\gamma^{l}$. The canonical expression for $\gamma^{l}$ given in Eq. (16a) and Eq. (16b) of Ref. 8 implicitly uses Eq. (9), which is valid only for the non-relativistic radial solutions $u_{l}$ and $\dot{u}_{l}$, Sec. II.2; (3) for the calculation of the SOC constants for the semicore $6p$-states we have used the $P_{\kappa=-2}$ large component of the Dirac solution for the $6p_{3/2}$ states, which, as demonstrated in Sec. II.3, gives better approximation for the actual energy splittings. For $6d$ and $5f$ levels, the SOC constants are calculated with the averaged components $P_{\ell}^{av}$, which describe the SO energy splittings of these states adequately; (4) we have taken into account the additional electron density ($Q$-density) close to the nuclear region by including the contributions from the small components of the $6p_{1/2}$ and $6p_{3/2}$ and other states.
Based on our calculations, the discussed difference in the treatment of relativistic effects can result in uncertainties up to 0.15 Å for lattice constants and to 26 GPa for bulk moduli even within the same chosen DFT functional (LDA, PBE or PBEsol), Tables 69.
@@
Our calculations include the full scale account of the SO coupling for valence states, which differs from other studies where the SO coupling is omitted or treated as a second variation step with certain approximations [24, 25, 10]. The calculated bulk moduli for Ac are in good correspondence with the experimental estimation of $B=24.5$ GPa ($0.25\times 10^{6}$ kg/cm2) made in Ref. [46] and the only theoretical value $B=25.9$ GPa, obtained on the basis of a tight-binding analysis in Ref. [48]. As for the calculation of the band structure of UO2, it is worth noting that in the literature it is often wrongly considered as having metallic band structure. Our calculations of UO2 with the SO coupling clearly demonstrate that it has a small gap ($\sim 0.2-0.4$ eV) at the Fermi energy, Fig. 5. Although the upper $5f$ occupied band and the next valence $5f$ band in UO2 slightly overlap, the situation is completely analogous to the calculated band structure of germanium, [50], which is a well known dielectric.

### v2

@@ Abstract
We examine the influence of the relativistic effects within the linear augmented plane wave method (LAPW) with the potential of general shape for solids and suggest a few ways to account them more accurately: (1) we introduce new radial dependencies for LAPW (Bloch-type) basis functions, based on two actual radial solutions of the Dirac equation for j=l1/2j=l-1/2 and $j=l+1/2$ one-electron states; (2) the canonical LAPW matrix elements for the spherically symmetric component of the potential, assuming non-relativistic radial wave functions, should be corrected; (3) we argue that for a realistic spin-orbit energy splitting of the semicore $6p$-states the spin-orbit interaction constant $\zeta(p)$ should be calculated with the $6p_{3/2}$ radial component; (4) in cases when two $j=l\pm 1/2$ components are occupied (for example for the $6p$ states of actinides) the electron density, associated with the small components of valence electrons, can be taken into the calculation scheme. We demonstrate that the new treatment for the relativistic effects is capable to change the equilibrium lattice constant up to 0.15 Å and the bulk modulus up to 26 GPa. We find that the electron density of valence electrons at the nucleus increases by 2.34.32.3-4.3 times due to the inclusion of small components, which can be essential for precise description of the potential and density close to the nuclear region, important for nuclear spectroscopies. In contrast to the common believe that in plain band structure treatment UO2 is a metal, we show that in the presence of the spin-orbit coupling UO2 has a small gap of forbidden states (0.20.40.2-0.4 eV) at the Fermi level, where the highest occupied and the lowest unoccupied $5f$ bands slightly overlap, as in calculations of the conduction and valence band in solid Ge.
@@ I Introduction
The aim of the present study is to increase the accuracy of the relativistic effects within the full electron full potential LAPW method (FLAPW) method for heavy elements as much as possible, while keeping the general scheme of the method unchanged. For that purpose (1) we introduce new basis functions, obtained from two independent solutions of the Dirac equation. In modern LAPW method one deals with two types of basis functions: Bloch-like LAPW basis functions and the atomic-like local orbitals (LO), which can be chosen to be fully relativistic (as e.g. the $p_{1/2}$ mentioned earlier [24, 25]). In our work the new type of radial basis functions concerns the Bloch-type LAPW basis functions. We do not consider here the choice of local orbitals, which is a separate problem [11, 26, 27]. Further, (2) we reconsider the matrix elements of the method, explicitly avoiding the use of hidden non-relativistic relations; (3) we correct the calculation of the spin-orbit (SO) coupling constant for $6p$ semicore states, based on the comparison the energy splittings between $6p_{1/2}$ and $6p_{3/2}$ components. In the present study we do not apply the second variation procedure for the SO-coupling. We use the direct treatment of the SO coupling in the full LAPW basis set thereby avoiding the approximations associated with the second variation step [10]. In this respect it is worth noting that in some relativistic calculations the authors tend to avoid the inclusion of the SO interaction [19, 20, 5], which, as discussed in Sec. II.3 below, can be connected with the overestimated values of SO coupling [6]. Finally, (4) we include the small components $Q$ in the calculations, which play an important role in the neighborhood of the nuclear region. These effects are considered and discussed in Sec. II.1, Sec. II.2, Sec. II.3, and Sec. II.4, correspondingly. In Sec. III we briefly review the results of our calculations for Ac, Th, ThO2 and UO2, and finally in Sec. IVwe discuss main conclusions and findings of our work. In our study we use various variants of the DFT functionals (see Sec. III below), which allows us to test the accuracy of the calculations.
@@ II.3 Calculated spin-orbit coupling constants, special treatment for $6p$-states
From the table 5 we can conclude that the actual energy splittings $\triangle E(p)$ between the $p_{1/2}$ and $p_{3/2}$ states are better approximated by the SO coupling constant $\zeta(p_{3/2})$ calculated using a single radial component $p_{3/2}$, for which $\triangle_{SO}^{*}(p)=3\zeta(p_{3/2})/2$. For example, for Ac we obtain $\triangle_{SO}^{*}(p)=6.09$ eV, for Th $\triangle_{SO}^{*}(p)=7.12$ eV etc. Although the values $\triangle_{SO}^{*}(p)$ are slightly smaller than the real energy differences $\triangle E(p)$, they approximate $\triangle E(p)$ better (i.e. 9119-11% vs 172717-27%) than the KH radial functions $P_{l=1}^{KH}$. The overestimation holds also for the solid case. Our data indicate that for example for the canonical (KH) Th variant the SO splitting between the $6p$-subband midpoints is 10.7 eV (with the with the minimum and maximum difference being 8.9 and 12.1 eV) whereas in the atomic case it is only 7.8 eV and the effective potential in the nuclear region is expected to remain basically unchanged. Therefore, in the following for a more realistic description of the SO splittings for the $6p$ semicore band states within the LAPW method we suggest to use only $p_{3/2}$ radial component. Since the actual equations for the SO coupling in LAPW (see e.g. Eq. (2) in [30]) differ from Eq. (14) and Eq. (15), and require three SO coupling constants $\zeta(p)$, $\dot{\zeta}(p)$, $\ddot{\zeta}(p)$, computed with two functions $P_{l=1}$ and $\dot{P}_{l=1}$, this implies that for their calculations we use $P_{3/2}$ (the $6p_{3/2}$ large component) and $\dot{P}_{3/2}$, which is the first energy derivative of the $6p_{3/2}$ large component.
For the other LAPW calculations of the SO interaction for the $d$ and $f$ valence states we use the standard procedure with the averaged radial components $P_{l}^{av}$ and $\dot{P}_{l}^{av}$, Eq. (3), because, as discussed above, they give good approximations of the energy splittings in the atomic case (i.e. $\triangle_{SO}(d)$ and $\triangle_{SO}(f)$). Our results obtained using this SO treatment in the LAPW method are presented in the tables 69 below as avD variants. The KH variants in the Tables correspond to the canonical SO treatment overestimating the $6p$ splitting as discussed earlier. Although in this Section we have considered the results with the PBE variant of DFT, the same conclusions can be drawn for other DFT functionals.
@@ III Application to actinides
The technical parameters of numerical calculations were as follows. For fcc Ac and fcc Th, in most cases the number of augmented plane waves was 137 and 274 (with SO), with $R_{MT}K_{j}\sim 10$, where $\vec{K}_{j}$ is the maximal wave vector, and $R_{MT}$ is the radius of the MT-sphere. For CaF2 cubic structures, the basis sets were 307 and 614 (with SO) for ThO2 ($R_{MT}K_{j}\sim 8.7$), 387 and 774 (with SO) for UO2 ($R_{MT}K_{j}\sim 9.5$). For the KH (canonical) basis two sets of $6p$ local functions ($u$ and $\dot{u}$) as described in Ref. [11]. Therefore, for the KH-calculations the total number of basis functions is 143 and 286 (with SO) for Ac and Th, 313 and 626 (with SO) for ThO2, and 393 and 786 for UO2. The following radii of MT-spheres have been used: for Ac 3.46 a.u. for all DFT functionals (i.e. LDA, PBEsol, PBE); for Th 3.05 a.u. (LDA), 3.12 a.u. (PBEsol and PBE); for ThO2 2.25 a.u. for both Th and O for all DFT functionals; for UO2 2.138 a.u. (LDA), 2.168 a.u. (PBEsol, PBE) for both U and O. It is worth noticing that within the chosen DFT functional we used the same $R_{MT}$ for all variants of the different radial functions (i.e. avD, avD(Q), KH with and without SOC). The dependence of the results of the new treatment on the chosen $R_{MT}$ is briefly discussed in Appendix C. The linear energy parameter $E_{\ell=1}$ for $p$-states was taken in the midpoint of the semicore $6p$ bands. The maximal number of $k$-points in the irreducible part (IP) of the Brillouin zone (BZ) for elemental actinides was 1505 ($\sim 70000$ for the whole BZ). For ThO2 and UO2 we have used a set of 240 $k$-points in IP of BZ ($\sim 11500$ for the whole BZ). The maximal value of the LAPW plane-wave expansion and the non-spherical density decomposition was $L_{max}=8$. We have taken into account the finite size of nuclei and used the tetrahedron method for the linear interpolation of energy between $k$-points [31]. The number of radial points inside the MT sphere region was increased to 4000-4200 for actinides and 900-1000 for oxygen. The enlarged number of radial points is analogous to the increase of the quality of the basis set, and this is certainly required for the actinides with 80 core electrons, which is a very large quantity. The calculations with the KH basis sets have been performed with additional localized atomic $p$-basis functions as described in Ref. [11]. We cannot perform the canonical (KH) calculations without supplying the basis set with local orbitals – such calculations are plagued with ghost states, or they crush and do not converge. In the case of new radial basis functions $P_{\ell}^{av}$ (avD) the calculations could proceed regularly until the convergence is reached. For that reason, all calculations marked in the Tables as avD or avD(Q) (with and without SO) have been carried out without local functions.
@@
The results of our calculations are listed in Table 6 for fcc Ac, Table 7 for fcc Th, Table 8 for cubic ThO2, and Table 9 for cubic UO2. ThO2 and UO2 are crystallized in the CaF2 structure. As discussed in the Introduction, most calculations of actinides [2, 4] and their oxides [5, 3, 35, 36] have been performed with the full potential LMTO method. Although nowadays the full potential LAPW (FLAPW) study of actinides [6, 7] and dioxides [37, 38] are also available, the new feature of the present FLAPW calculations is the complete treatment of the SO couplings. Recently, high precision PBE calculations of actinides and their oxides without SO coupling were reported in Ref. [39]. They can be compared with our results for the PBE variant with the canonical (the KH basis) calculation without SO coupling. Our equilibrium lattice constants for Ac and Th agree with those given in the Supplementary information of [39] within 0.2% and 0.3%, respectively, while for the dioxides (ThO2, UO2), the difference is 1.2%. This discrepancy is explained by the use of different technical parameters (MT radii, etc.), local orbitals [11], and even fitting functions (we use the Murnaghan equation of state).
It is worth mentioning that there are several theoretical studies of UO2 with correlation effects (Hubbard repulsion) [40, 41, 42, 43, 44, 45, 46]. Such an approach however lies beyond the scope of the present work, which focuses on the peculiarities of the inclusion of relativistic effects.
@@
It is also worth mentioning that in contrast to Th, our DFT calculations of the fcc structure of Ac appreciably overestimate its lattice constant even for LDA. This however was also noticed e.g. in Ref. [51] ($a=5.503$ Å in LDA) and very recently was also confirmed for the PBE LAPW variant without SOC [39] ($a=5.669$ Å). Therefore, the effect should be attributed to the peculiarity of the band structure of this element. Our calculations indicate that this feature becomes more pronounced with increasing quality of the basis set. For example, decreasing the non-spherical components of electron density to $L_{max}=6$ in the avD(Q) basis leads to a smaller lattice constants: 5.536 Å ($\delta a=-0.037$ Å) in LDA, 5.736 Å (0.018-0.018 Å) in PBE, 5.597 Å (0.032-0.032 Å) in PBEsol. Further, decrease of the basis set to only 65 functions results in 5.500 Å (total $\delta a=-0.073$ Å in LDA) 5.608 Å (0.146-0.146 Å) in PBE, 5.539 Å (0.09-0.09 Å) in PBEsol, compare with Table 6. Possibly, some properties of the phonon spectrum and mean square displacements of atoms in solid Ac make the description using poor basis sets more adequate to the experimental data. Owing to its scarcity and radioactivity, the experimental bulk modulus of Ac is unknown. To the best of our knowledge in the literature there is only an estimated (not directly measured) value of 24.5 GPa, listed in Ref. [49]. Theoretical bulk moduli are $B=25.9$ GPa, obtained on the basis of a tight-binding (LDA) analysis [51], and $B=23.9$ GPa in the canonical PBE LAPW variant without SOC [39]. All these values are in good correspondence with our data, Table 6.
@@
Figure 5: The upper panel of the calculated band structure of UO2 with the spin-orbit coupling, PBE calculation. The highest occupied $5f$ electron band is shown by solid line, lowest unoccupied $5f$ bands by dashed lines. The vertical gap with $\triangle E$ from 0.20.40.2-0.4 eV is visible. (The zero energy here corresponds to the position of the first $5f$ band at the $\Gamma$ point, $E_{F}=-0.16$ eV.)
@@ IV Conclusions
Within the LAPW method we have presented a few ways to include the relativistic effects more completely and consistently: (1) we have used new radial functions $P_{\ell}^{av}$ and $\dot{P}_{\ell}^{av}$ for the Bloch-type basis states. The functions are obtained by finding the large components $P_{\kappa=\ell}$, $P_{\kappa=-\ell-1}$ of the Dirac solutions independently for the $j=\ell-1/2$ and $j=\ell+1/2$ states and then averaging them explicitly by means of Eq. (3). Although we have found that new radial basis functions bring more stability to the self-consistent-field procedure, they can be enriched with local atomic-like orbitals [24, 25, 11, 26, 27] as done with the canonical (KH) basis set; (2) we have corrected the LAPW expressions for $a_{l}^{n}$, $b_{l}^{n}$, Eq. (11a), Eq. (11b), and for the matrix elements in the spherical ($L=0$) component of the total potential, using Eq. (13) for $\gamma^{l}$. The canonical expression for $\gamma^{l}$ given in Eq. (16a) and Eq. (16b) of Ref. 8 implicitly uses Eq. (9), which is valid only for the non-relativistic radial solutions $u_{l}$ and $\dot{u}_{l}$, Sec. II.2; (3) for the calculation of the SOC constants $\zeta(p)$, $\dot{\zeta}(p)$, $\ddot{\zeta}(p)$ for the semicore $6p$-states we have used the large component of the Dirac solution for the $6p_{3/2}$ states $P_{\kappa=-2}$ and its energy derivative $\dot{P}_{\kappa=-2}$, which, as demonstrated in Sec. II.3 for the atomic state and in our band structure calculations, gives better approximation for the actual energy splittings. For $6d$, $5f$ and high $\ell$ levels, the SOC constants are calculated with the new averaged radial components $P_{\ell}^{av}$, $\dot{P}_{\ell}^{av}$, because they describe the SO energy splittings adequately (i.e. here $P_{\ell}^{av}$, $\dot{P}_{\ell}^{av}$ simply substitute the canonical radial functions $P_{\ell}^{KH}$, $\dot{P}_{\ell}^{KH}$ in the LAPW expressions for SOC); (4) we have taken into account the additional electron density ($Q$-density) close to the nuclear region by including the contributions from the small components of the $6p_{1/2}$ and $6p_{3/2}$ and other states.
Our calculations include the full scale account of the SO coupling for valence states, which differs from other studies where the SO coupling is omitted or treated as a second variation step with certain approximations [24, 25, 10]. Based on our study, the discussed difference in the treatment of relativistic effects can result in uncertainties up to 0.15 Å for lattice constants and to 26 GPa for bulk moduli even within the same chosen DFT functional (LDA, PBE or PBEsol), Tables 69.
@@
As for the calculation of the band structure of UO2, it is worth noting that in the literature it is often wrongly considered as having metallic band structure. Our study of UO2 with the full treatment of the SO coupling clearly demonstrates that it has a small gap ($\sim 0.2-0.4$ eV) at the Fermi energy, Fig. 5. Although the upper $5f$ occupied band and the next valence $5f$ band in UO2 slightly overlap, the situation is completely analogous to the calculated band structure of germanium [53]. Since in reality germanium is a dielectric, the same effect can occur also in UO2.